OpenStudy (anonymous):

two integers differ by 12 and the sum of their squares is 74. find the integers ???? :(

OpenStudy (anonymous):

x = integer 1 y = integer 2 y - x = 12 y^2 + x^2 = 74 now you have a system of equations solve the 1st equation for x, in terms of y substitute that into the 2nd equation and solve for y

OpenStudy (campbell_st):

well let 1 integer = x the other is x + 12 then \[x^2 + (x + 12)^2 = 74\] so simplifying \[2x^2 + 24x + 144 = 74\] or \[x^2 + 12x + 35 = 0\] fyou can solve the quadratic by factoring....

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait but isn't the integer ment to "differ" so it's - instead of +

OpenStudy (anonymous):

x differrs from x+ 12 by 12 units this is the same as (x+12) - 12 = x

OpenStudy (anonymous):

x + 12 = y y - 12 = x

OpenStudy (anonymous):

sorry I still don't get it because....this is what I got (x^2-12)=74

OpenStudy (anonymous):

mistake (x-12)^2=74

OpenStudy (anonymous):

makes no difference

OpenStudy (anonymous):

then I simplified it and got x^2 - 24x + 144=74 = x^2 - 24x + 70=0

OpenStudy (anonymous):

it is the sum of the squares either \[x^2+(x-12)^2=74\] or \[x^2+(x+12)^2=74\]

OpenStudy (anonymous):

thank you :)